# The space of immersed polygons

### v1

@@ Abstract.
We use the Schwarz–Christoffel formula to show that for every $n\geq 3$, the space of labelled immersed $n$-gons in the plane up to similarity is homeomorphic to ${\mathbb{R}}^{2n-4}$. It follows that the space of labelled simple $n$-gons up to similarity is homeomorphic to ${\mathbb{R}}^{2n-4}$ if $n\in\{3,4,5\}$, which confirms one more case of a conjecture of González and Sedano-Mendoza.
@@ 1. Introduction
For $n=3$, this result is well-known (see e.g. [Bae24]) and elementary. Here is a proof: the map that sends a triangle with counter-clockwise vertices $w_{1}$, $w_{2}$, $w_{3}$ to $(w_{3}-w_{1})/(w_{2}-w_{1})$ induces a homeomorphism from $\operatorname{LSP}(3)/\operatorname{Aff}({\mathbb{C}})$ to the upper half-plane ${\mathbb{H}}$, which is homeomorphic to ${\mathbb{R}}^{2}$ via the map $x+iy\mapsto x+i\log y$. For $n=4$, the result was first proved in [GLL16] and then again in [GSM23] with different methods. In [GSM23], González and Sedano-Mendoza conjectured that the result holds for every $n\geq 5$ as well. The above corollary settles the case $n=5$ and reproves the previous cases in a unified way. For general $n\geq 3$, it is known that $\operatorname{LSP}(n)/\operatorname{Aff}(n)$ is simply-connected [Gon18, Theorem 3.2].
@@
The works [CDR03] and [SW14] in particular study simple polygons with fixed side lengths. Some of their results can be restated as follows. Given positive real numbers $\ell_{1},\ldots,\ell_{n}$ such that $\ell_{j}<\sum_{k\neq j}\ell_{k}$ for every $j$, the subset $X(\ell)$ of polygons in $\operatorname{LSP}(n)/\operatorname{Aff}(n)$ whose vector of side lengths is a constant multiple of $\ell=(\ell_{1},\ldots,\ell_{n})$ is contractible [CDR03, Corollary 3]. If furthermore the $\ell_{j}$ cannot partitioned into two sets with the same sum, then $X(\ell)$ is homeomorphic to ${\mathbb{R}}^{n-3}$ [SW14, Theorem 7]. If the additionnal hypothesis on subsets with equal sums could be removed in this result of Shimamoto–Wooters, then the conjecture of González and Sedano-Mendoza would follow, since the set of realizable vectors of side lengths up to scaling is easily seen to be homeomorphic to ${\mathbb{R}}^{n-1}$.
@@ 2. Schwarz–Christoffel tranformations
For any choice of parameters as above, the Schwarz–Christoffel transformation $F$ is a local homeomorphism, holomorphic in ${\mathbb{H}}$, whose restriction to ${}{\mathbb{R}}\cup\{\infty\}$ is an immersed $n$-gon in ${\mathbb{C}}$ with interior angle $\alpha_{j}\pi$ at vertex $F(z_{j})$ for every $j\in\{1,\ldots,n\}$. The resulting polygon is simple if and only if $F$ is injective. Conversely, any immersed $n$-gon in ${\mathbb{C}}$ arises as the restriction to ${}{\mathbb{R}}\cup\{\infty\}$ of a Schwarz–Christoffel transformation, unique up to pre-composition with a map of the form $z\mapsto az+b$ with $a>0$ and $b\in{\mathbb{R}}$.
@@ 4. Embedded polygons
Let $\phi$ be an immersed $n$-gon which is not simple. We can perturb $\phi$ to a nearby non-simple immersed polygon $\psi$ where all self-intersections happen away from the vertices and are transverse. Since $\psi$ is not simple, there is some point $p$ in the complement of its image such that the winding number of $\psi$ around $p$ is at least $2$ (winding numbers are non-negative for immersed polygons).
@@
Suppose that $n=5$. Then each side of $\psi$ must be oriented to the left with respect to $p$, otherwise the winding number would be strictly less than $2=4\cdot\frac{1}{2}$ due to the negative contributions. Let us form form triangles with $p$ and each side of $\psi$. Then the sum of the angles $\gamma_{k}$ of these triangles at the point $p$ is equal to $2\pi k$ where $k\geq 2$ is the winding number of $\psi$ around $p$. On the other hand, by adding all the interior angles of these triangles, we find that

### v2

@@ Abstract.
We use the Schwarz–Christoffel formula to show that for every $n\geq 3$, the space of labelled immersed $n$-gons in the plane up to similarity is homeomorphic to ${\mathbb{R}}^{2n-4}$. We then prove that all immersed triangles, quadrilaterals, and pentagons are embedded, from which it follows that the space of labelled simple $n$-gons up to similarity is homeomorphic to ${\mathbb{R}}^{2n-4}$ if $n\in\{3,4,5\}$. This was first shown by Gonzáles and López-López for $n=4$ and conjectured to be true for every $n\geq 5$ by González and Sedano-Mendoza.
@@ 1. Introduction
For $n=3$, this result is well-known (see e.g. [Bae24]) and elementary. Here is a proof: the map that sends a triangle with counter-clockwise vertices $w_{1}$, $w_{2}$, $w_{3}$ to
$\frac{w_{3}-w_{1}}{w_{2}-w_{1}}$
induces a homeomorphism from $\operatorname{LSP}(3)/\operatorname{Aff}({\mathbb{C}})$ to the upper half-plane ${\mathbb{H}}$, which is homeomorphic to ${\mathbb{R}}^{2}$ via the map $x+iy\mapsto(x,\log y)$. For $n=4$, the result was first proved in [GLL16] and then again in [GSM23] with different methods. In [GSM23], González and Sedano-Mendoza conjectured that the result holds for every $n\geq 5$ as well. The above corollary settles the case $n=5$ and reproves the previous cases in a unified way. For general $n\geq 3$, it is known that $\operatorname{LSP}(n)/\operatorname{Aff}({\mathbb{C}})$ is simply-connected [Gon18, Theorem 3.2].
@@
The works [CDR03] and [SW14] in particular study simple polygons with fixed side lengths. Some of their results can be restated as follows. Given positive real numbers $\ell_{1},\ldots,\ell_{n}$ such that $\ell_{j}<\sum_{k\neq j}\ell_{k}$ for every $j$, the subset $X(\ell)$ of polygons in $\operatorname{LSP}(n)/\operatorname{Aff}({\mathbb{C}})$ whose vector of side lengths is a constant multiple of $\ell=(\ell_{1},\ldots,\ell_{n})$ is contractible [CDR03, Corollary 3]. If furthermore the $\ell_{j}$ cannot partitioned into two sets with the same sum, then $X(\ell)$ is homeomorphic to ${\mathbb{R}}^{n-3}$ [SW14, Theorem 7]. If the additionnal hypothesis on subsets with equal sums could be removed in this result of Shimamoto–Wooters, then the conjecture of González and Sedano-Mendoza would follow, since the set of realizable vectors of side lengths up to scaling is homeomorphic to ${\mathbb{R}}^{n-1}$, and a fiber bundle over a contractible base is always trivial.
This is all to say that the statement of Corollary 1.2 is likely to be true for every $n\geq 3$ even though that cannot be deduced from Theorem 1.1.
@@ 2. Schwarz–Christoffel tranformations
For any choice of parameters as above, the Schwarz–Christoffel transformation $F$ is a local homeomorphism, holomorphic in ${\mathbb{H}}$, whose restriction to ${}{\mathbb{R}}\cup\{\infty\}$ is an immersed $n$-gon in ${\mathbb{C}}$ with interior angle $\alpha_{j}\pi$ at vertex $F(z_{j})$ for every $j\in\{1,\ldots,n\}$. The resulting polygon is simple if and only if $F$ is injective. Conversely, any immersed $n$-gon in ${\mathbb{C}}$ arises as the restriction to ${}{\mathbb{R}}\cup\{\infty\}$ of a Schwarz–Christoffel transformation, unique up to pre-composition by an affine map $z\mapsto az+b$ with $a>0$ and $b\in{\mathbb{R}}$.
@@ 4. Embedded polygons
Let $\phi$ be an immersed $n$-gon which is not simple. We can perturb $\phi$ to a nearby non-simple immersed polygon $\psi$ where all self-intersections happen away from the vertices and are transverse. Since $\psi$ is not simple, there is some point $p$ in the complement of its image such that the winding number of $\psi$ around $p$ is at least $2$. Indeed, winding numbers are non-negative for immersed polygons and the winding number always takes 3 distinct values around a transverse intersection point. We can further assume that $p$ does not lie on any of the lines that contain the sides of $\psi$.
@@
Suppose that $n=5$. Then each side of $\psi$ must be oriented to the left with respect to $p$, otherwise the winding number would be strictly less than $2=4\cdot\frac{1}{2}$ due to the negative contributions. Let us form triangles with $p$ and each side of $\psi$. Then the sum of the angles $\gamma_{k}$ of these triangles at the point $p$ is equal to $2\pi k$ where $k\geq 2$ is the winding number of $\psi$ around $p$. On the other hand, by adding all the interior angles of these triangles, we find that